The perimeter of a rectangle is 968 cm. The width is one-third as long as the length. Find the dimensions.
Answer by fractalier(6550)

Call the length and width, L and W.
We have W = (1/3)L or L = 3W.
Perimeter P = 2L + 2W = 968
Substituting the first equation into the second gives
2(3W) + 2W = 968
6W + 2W = 968
8W = 968
W = 121 cm
and since this is one-third of the length,
L = 363 cm